a quadrilateral has four sides. Quadrlateral ABCD has three sides that have equal measures. The measure of the fourth side is 8cm longer than each of teh over sides. If the perimeter of the quadriladeral is 28m find the measure of each side using the correct number of significant digits
Answer by oberobic(2304)

Three sides are equal:
A = B = C
.
The other side is 8 cm longer
D = A+8 = B+8 = C+8
.
Perimeter = 28
.
3A +A+8 = 28
4A +8 = 28
4A = 20
A = 5
so
B = 5
and
C = 5
.
D = 5+8 = 13
.
15+13 = 28
Correct.
.
The sides are 5, 5, 5, and 13 cm.
.
Done.